# Learning Robust and Correct Controllers Guided by Feasibility-Aware Signal Temporal Logic via BarrierNet

### v2

@@ IV-B HOCBFs for STL specifications
The authors of [lindemann2018control] introduced the use of time-varying CBFs to enforce the satisfaction of STL specifications. However, [lindemann2018control] considers only CBFs of relative degree one, and the construction of these CBFs is illustrated through examples rather than provided in a general or systematic form. In this work, we extend their approach to high-order CBFs and develop a general, algorithmic procedure for constructing such HOCBFs. Moreover, we make these HOCBFs trainable, eliminating the need for manual design and enabling the resulting controller to further improve its performance through training.
@@
  • Category I: predicates that are already satisfied at $t=0$ and whose enclosing temporal operator also begins at $t=0$. For example, $\mu_{1}$ in $G_{[0,5]}\mu_{1}$ and in $G_{[0,5]}\mu_{1}\land\mu_{2}$, where $h_{1}(\mathbf{x}_{0})\geq 0$. Such predicates typically encode safety requirements, such as obstacle avoidance in robotic applications.
  • Category II: predicates enclosed by $F_{[t_{a},t_{b}]}$ that do not belong to Category I, e.g., $\mu_{1}$ in $F_{[2,5]}\mu_{1}$ and $\mu_{2}$ in $F_{[0,5]}\mu_{2}\land\mu_{3}$, where $h_{2}(\mathbf{x}_{0})<0$.
  • Category III: predicates enclosed by $G_{[t_{a},t_{b}]}$ that do not belong to Category I. For example, $\mu_{1}$ in $G_{[2,5]}\mu_{1}\land\mu_{2}$. Note that Assumption 1 rules out formulae such as $G_{[0,5]}\mu_{1}$ with $h_{1}(\mathbf{x}_{0})<0$, which are infeasible from the initial condition.
@@
Constraint (22a) ensures that the HOCBF is positive at the initial time, i.e., $b(\mathbf{x}_{0},0)>0$. Constraint (22b) ensures that $h(\mathbf{x})\geq b(\mathbf{x},t)$ for some interval of time before $t_{b}$. Under (22a) and (22b), the forward invariance of the superlevel set of $b(\mathbf{x},t)$ enforces the satisfaction of $F_{[t_{a},t_{b}]}\mu$. The third constraint (22c) ensures that the superlevel set of $b(\mathbf{x},t)$ is nonempty for $t<t_{a}$. As discussed later, we delete this HOCBF once $h(\mathbf{x})>0$ for some $t\geq t_{a}$, so we do not require the superlevel set of $b(\mathbf{x},t)$ to remain nonempty thereafter.
@@
The difference is that (24b) ensures that $h(\mathbf{x})\geq b(\mathbf{x},t)$ over some interval prior to $t_{a}$, so that, combined with the forward invariance of the superlevel set of $b(\mathbf{x},t)$, the requirement $G_{[t_{a},t_{b}]}\mu$ is satisfied. When $c>0$ is small enough, the superlevel set of $b(\mathbf{x},t)$ is always nonempty under Assumption 1. We choose the exponential function (23) for always instead of a linear function (21) because it satisfies:
@@
Consider $\varphi=F_{[0,2]}\mu_{1}\land F_{[2,4]}\mu_{2}$, where $h_{1}(\mathbf{x})=R_{1}-\sigma(l(\mathbf{x})-\mathbf{o}_{1}),h_{2}(\mathbf{x})=R_{2}-\sigma(l(\mathbf{x})-\mathbf{o}_{2}).$ At time $t$, the invariant sets induced by their HOCBFs are $\mathcal{B}^{\mathrm{av}}\big(\mathbf{o}_{1},R_{1}+\gamma_{1}(t)\big),\mathcal{B}^{\mathrm{av}}\big(\mathbf{o}_{2},R_{2}+\gamma_{2}(t)\big).$ To avoid conflicts between these two HOCBFs on the interval $[0,2]$, we: (1) enforce that their sets have a nonempty intersection on $[0,2]$; and (2) delete $b_{1}$ once $h_{1}(\mathbf{x})>0$. A sufficient condition for (1) is $\mathcal{B}^{\mathrm{av}}\left(\mathbf{o}_{1},R_{1}+\gamma_{1}(2)\right)\cap\mathcal{B}^{\mathrm{av}}\left(\mathbf{o}_{2},R_{2}+\gamma_{2}(2)\right)\neq\emptyset.$ Using the triangle inequality of $\rho$, this holds if $\gamma_{2}(2)\;\geq\;\sigma(\mathbf{o}_{1}-\mathbf{o}_{2})-\gamma_{1}(2)-R_{1}-R_{2}.$
Deletion rules. For each predicate wrapped with $G_{[t_{a},t_{b}]}\mu$, we delete the associated HOCBF at $t=t_{b}$. For $F_{[t_{a},t_{b}]}\mu$, we delete the HOCBF once $h(\mathbf{x})>0$ for some $t\geq t_{a}$. For $F_{[t_{a},t_{b}]}\bigwedge_{j=1}^{N}\mu_{j}$, all corresponding HOCBFs are deleted together once $h_{j}(\mathbf{x})>0$ for all $j$. Note that $G_{[t_{a},t_{b}]}(\mu_{1}\land\mu_{2})$ is equivalent to $G_{[t_{a},t_{b}]}\mu_{1}\land G_{[t_{a},t_{b}]}\mu_{2}$, but $F_{[t_{a},t_{b}]}(\mu_{1}\land\mu_{2})$ is not equivalent to $F_{[t_{a},t_{b}]}\mu_{1}\land F_{[t_{a},t_{b}]}\mu_{2}$ since the latter allows asynchronous satisfaction.
@@
where $\bm{\omega}$ is omitted, $h_{k}(\mathbf{x})=s_{k}\big(R_{k}-\sigma(l(\mathbf{x})-\mathbf{o}_{k})\big)$, $D_{j,k}=\min(s_{j}+s_{k},0)\times\inf,$ so that when $s_{j}=s_{k}=-1$ (both avoidance tasks), the condition is released. For predicates in Category I, we simply set $\gamma(t)=0$. Geometrically, the conditions in (27) ensure that, at each predicate’s ending time, the feasible sets of all previously active HOCBFs remain mutually compatible. To illustrate the idea, consider two predicates $\mu_{1}$ and $\mu_{2}$ with ending times $t_{b}^{1}\leq t_{b}^{2}$. Since $b_{1}$ must be maintained until $t_{b}^{1}$, we require that at $t_{b}^{1}$ the feasible region induced by $b_{2}$ does not conflict with that of $b_{1}$, then the corresponding compatibility condition becomes:
@@
Because $\gamma(t)$ in (21) and (23) is non-increasing, the temporal operators are enforced as follows:
@@
  • For $F_{[t_{a},t_{b}]}\bigwedge_{j=1}^{N}\mu_{j}$, constraint (22b) ensures the existence of a common $t^{\prime}\in[t_{a},t_{b}]$ where all $h_{j}(\mathbf{x}(t^{\prime}))\geq 0$.
@@
Consider system (1) satisfying Assumptions 1 and 2, and HOCBFs constructed by (19) and (20) that satisfy (22), (24), and (27). Suppose each HOCBF satisfies $\psi_{i,j}(\mathbf{x}_{0},0,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})\geq 0$ for $i=0,\ldots,m-1$$j=1,\ldots,M$. Then any control input $\mathbf{u}(t)$ that satisfies (7) for all undeleted HOCBFs guarantees satisfaction of the STL specification $\varphi$.
@@
Constraints (22a) and (24a) ensure $b_{j}(\mathbf{x}_{0},0)\geq 0$ for all predicates. Since $\psi_{i,j}(\mathbf{x}_{0},0,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})\geq 0$ for all $i$, Theorem 1 guarantees that any control input satisfying (7) yields $b_{j}(\mathbf{x},t)\geq 0$ for all undeleted HOCBFs. Because $\gamma_{j}(t)$ is non-increasing, (22b) ensures $\exists t^{\prime}\in[t_{a},t_{b}]$, $h(\mathbf{x}(t^{\prime}))\geq 0$ for $F{[t_{a},t_{b}]}\mu$, while (24b) ensures $\forall t^{\prime}\in[t_{a},t_{b}]$, $h(\mathbf{x}(t^{\prime}))\geq 0$ for $G{[t_{a},t_{b}]}\mu$. For a formula in the form of $F_{[t_{a},t_{b}]}\land_{j=1}^{N}\mu_{j}$, (22b) ensures $\exists t^{\prime}\in[t_{a},t_{b}]$, $h_{j}(\mathbf{x}(t^{\prime}))\geq 0$ for all $j$. Hence every predicate in $\varphi$ is satisfied over its required temporal interval, ensuring $(\mathbf{x},0)\models\varphi$. ∎
@@ IV-C Unified Robustness Metric
In Sec. IV-B we constructed either (i) a time-invariant HOCBF $b_{j}(\mathbf{x})$ (for Category I predicates), or (ii) a time-varying HOCBF $b_{j}(\mathbf{x},t)$ (for Category II and III predicates), together with a deletion rule that removes the corresponding constraint once the predicate is satisfied (for $F_{[t_{a},t_{b}]}\mu$ operators) or at the terminal time (for $G_{[t_{a},t_{b}]}\mu$ operators). Before deletion, each HOCBF imposes a highest-order constraint $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})\geq 0$, which must remain feasible for the QP. This subsection introduces a unified robustness combining: (i) feasibility robustness, and (ii) STL temporal robustness based on the exponential robustness in [liu2023robust], ensuring that the controller remains feasible while satisfying the STL specification.
@@
Feasibility robustness. As shown in Thm. 3, the QP is feasible at time $t$ whenever $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$. Thus the instantaneous feasibility margin is
$\rho^{\mathrm{fea}}_{j}(\mu,\mathbf{x},t):=\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p}).$ (31)
This margin must stay positive until the HOCBF is deleted. Let the deletion time be
$t_{\mathrm{del}}^{j}=\begin{cases}\inf\{t\geq t_{a}^{j}:h_{j}(\mathbf{x}(t))\geq 0\},&F_{[t_{a}^{j},t_{b}^{j}]}\mu\text{-type},\\[2.84526pt] t_{b}^{j},&G_{[t_{a}^{j},t_{b}^{j}]}\mu\text{-type}.\end{cases}$ (32)
Since the STL fragment considered in (4b) excludes the temporal until operator, thus the temporal feasibility requirement $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$ until deletion” is equivalent to the standard until-free semantics:
$\varphi_{j}^{\mathrm{fea}}:=F_{[t_{a}^{j},t_{b}^{j}]}\mathrm{del}_{j}\wedge G_{[0,t_{b}^{j}]}(\mathrm{del}_{j}\lor\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0),$ (33)
where $\mathrm{del}_{j}$ denotes the event of deletion. With the construction above, we now have $M$ predicates $\mu_{j}$, $j=1,\ldots,M$, spanning Categories I–III. Each predicate $\mu_{j}$ is associated with an STL subformula $\varphi_{j}^{\mathrm{fea}}$ that enforces QP feasibility until deletion, given by expression (33). Each $\varphi_{j}^{\mathrm{fea}}$ is itself an STL formula in the same fragment as (4b), built from atomic predicates of the form $\psi_{m-1,j}>0$. To capture the control bounds (10c), we introduce a single STL subformula which must hold for all $t\in[0,T]$:
$\varphi^{\mathcal{U}}:=G_{[0,T]}\big((\mathbf{u}(\mathbf{x}(t))-\mathbf{u}_{\min}\geq\mathbf{0})\wedge(\mathbf{u}_{\max}-\mathbf{u}(\mathbf{x}(t))\geq\mathbf{0})\big).$ (34)
@@
Since $\widetilde{\varphi}$ lies in the same STL fragment as (4b), its robustness can be evaluated using the exponential robustness in (28)–(30). We denote the resulting unified scalar robustness by
@@
Theorem 5 (Feasibility-Aware STL Guarantee).
Consider system (1) with HOCBFs constructed as in Sec. IV-B and the associated feasibility subformulas $\varphi_{j}^{\mathrm{fea}}$ in (33) and control bound subformula in (34). Assume that at every time $t$ the dQP (10) enforces the highest-order HOCBF constraints and deletion rules of Sec. IV-B. If
$\rho^{\mathrm{uni}}(\mathbf{x},0)=\rho(\widetilde{\varphi},\mathbf{x},0)\;\geq 0,$ (37)
then the closed-loop trajectory satisfies the original STL specification $\varphi$, and the dQP remains feasible for all $t$.
@@
By soundness of the exponential robustness in (28)–(30), $\rho^{\mathrm{uni}}(\mathbf{x},0)\geq 0$ is equivalent to $(\mathbf{x},0)\models\widetilde{\varphi}$. From the definition (35), this implies simultaneously $(\mathbf{x},0)\models\varphi$, $(\mathbf{x},0)\models\varphi^{\mathcal{U}}$ and $(\mathbf{x},0)\models\varphi_{j}^{\mathrm{fea}}$ for all $j=1,\ldots,M$. Satisfaction of each $\varphi_{j}^{\mathrm{fea}}$ means that: (i) the deletion event $\mathrm{del}_{j}$ occurs within $t_{a}^{j},t_{b}^{j}]$, and (ii) $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$ holds for all $t<t_{\mathrm{del}}^{j}$. By Thm. 3, this guarantees pointwise feasibility of the dQP (10) for all times while $b_{j}$ is active. Together with Thm. 2, which ensures forward invariance of the HOCBF-defined sets under feasible dQP solutions, we conclude that the QP remains feasible for all $t$ and that all predicates in $\varphi$ are satisfied over their required temporal intervals. Hence $(\mathbf{x},0)\models\varphi$. ∎

### v3

@@ IV-B HOCBFs for STL specifications
The authors of [16] introduced the use of time-varying CBFs to enforce the satisfaction of STL specifications. However, [16] considers only CBFs of relative degree one, and the construction of these CBFs is illustrated through examples rather than provided in a general or systematic form. In this work, we extend their approach to high-order CBFs and develop a general, algorithmic procedure for constructing such HOCBFs. Moreover, we make these HOCBFs trainable, eliminating the need for manual design and enabling the resulting controller to further improve its performance through training.
@@
  • Category I: predicates that are already satisfied at $t=0$ and whose enclosing temporal operator also begins at $t=0$. For example, $\mu_{1}$ in $G_{[0,5]}\mu_{1}$ and in $G_{[0,5]}(\mu_{1}\land\mu_{2})$, where $h_{1}(\mathbf{x}_{0})\geq 0$. Such predicates typically encode safety requirements, such as obstacle avoidance in robotic applications.
  • Category II: predicates enclosed by $F_{[t_{a},t_{b}]}$ that do not belong to Category I, e.g., $\mu_{1}$ in $F_{[2,5]}\mu_{1}$ and $\mu_{2}$ in $F_{[0,5]}(\mu_{2}\land\mu_{3})$, where $h_{2}(\mathbf{x}_{0})<0$.
  • Category III: predicates enclosed by $G_{[t_{a},t_{b}]}$ that do not belong to Category I. For example, $\mu_{1}$ in $G_{[2,5]}(\mu_{1}\land\mu_{2})$. Note that Assumption 1 rules out formulae such as $G_{[0,5]}\mu_{1}$ with $h_{1}(\mathbf{x}_{0})<0$, which are infeasible from the initial condition.
@@
Constraint (22a) ensures that the HOCBF is positive at the initial time, i.e., $b(\mathbf{x}_{0},0)>0$. Constraint (22b) ensures that $h(\mathbf{x})\geq b(\mathbf{x},t)$ before $t_{b}$. Under (22a) and (22b), the forward invariance of the superlevel set of $b(\mathbf{x},t)$ enforces the satisfaction of $F_{[t_{a},t_{b}]}\mu$. The third constraint (22c) ensures that the superlevel set of $b(\mathbf{x},t)$ is nonempty for $t<t_{a}$. As it will be discussed later, we delete this HOCBF once $h(\mathbf{x})\geq 0$ for some $t\geq t_{a}$, so we do not require the superlevel set of $b(\mathbf{x},t)$ to remain nonempty thereafter.
@@
The difference is that (24b) ensures $h(\mathbf{x})\geq b(\mathbf{x},t)$ before $t_{a}$, so that, combined with the forward invariance of the superlevel set of $b(\mathbf{x},t)$, the requirement $G_{[t_{a},t_{b}]}\mu$ is satisfied. When $c>0$ is small enough, the superlevel set of $b(\mathbf{x},t)$ is always nonempty under Assumption 1. We choose the exponential function (23) for always instead of a linear function (21) because it satisfies:
@@
Consider $\varphi=F_{[0,2]}\mu_{1}\land F_{[2,4]}\mu_{2}$, where $h_{1}(\mathbf{x})=R_{1}-\sigma(l(\mathbf{x})-\mathbf{o}_{1}),h_{2}(\mathbf{x})=R_{2}-\sigma(l(\mathbf{x})-\mathbf{o}_{2}).$ At time $t$, the invariant sets induced by their HOCBFs are $\mathcal{B}^{\mathrm{re}}\big(\mathbf{o}_{1},R_{1}+\gamma_{1}(t)\big),\mathcal{B}^{\mathrm{re}}\big(\mathbf{o}_{2},R_{2}+\gamma_{2}(t)\big).$ To avoid conflicts between these two HOCBFs on the interval $[0,2]$, we: (1) enforce that their invariant sets have a nonempty intersection on $[0,2]$; and (2) delete $b_{1}$ once $h_{1}(\mathbf{x})>0$. A sufficient condition for (1) is $\mathcal{B}^{\mathrm{re}}\left(\mathbf{o}_{1},R_{1}+\gamma_{1}(2)\right)\cap\mathcal{B}^{\mathrm{re}}\left(\mathbf{o}_{2},R_{2}+\gamma_{2}(2)\right)\neq\emptyset.$ Using the triangle inequality of $\sigma$, this holds if $\gamma_{2}(2)\;\geq\;\sigma(\mathbf{o}_{1}-\mathbf{o}_{2})-\gamma_{1}(2)-R_{1}-R_{2}.$
For each predicate wrapped with $G_{[t_{a},t_{b}]}\mu$, we delete the associated HOCBF at $t=t_{b}$. For $F_{[t_{a},t_{b}]}\mu$, we delete the HOCBF once $h(\mathbf{x})\geq 0$ for some $t\geq t_{a}$. For $F_{[t_{a},t_{b}]}(\bigwedge_{j=1}^{N}\mu_{j})$, all corresponding HOCBFs are deleted together once $h_{j}(\mathbf{x})\geq 0$ for all $j$. Note that $G_{[t_{a},t_{b}]}(\mu_{1}\land\mu_{2})$ is equivalent to $G_{[t_{a},t_{b}]}\mu_{1}\land G_{[t_{a},t_{b}]}\mu_{2}$, but $F_{[t_{a},t_{b}]}(\mu_{1}\land\mu_{2})$ is not equivalent to $F_{[t_{a},t_{b}]}\mu_{1}\land F_{[t_{a},t_{b}]}\mu_{2}$ since the latter allows asynchronous satisfaction.
@@
where $\bm{\omega}$ is omitted, $h_{k}(\mathbf{x})=s_{k}\big(R_{k}-\sigma(l(\mathbf{x})-\mathbf{o}_{k})\big)$, $D_{j,k}=\min(s_{j}+s_{k},0)\times\inf,$ so that when $s_{j}=s_{k}=-1$ (both avoidance tasks), the condition is released. For predicates in Category I, we simply set $\gamma(t,\bm{\omega})=0$. Geometrically, the conditions in (27) ensure that, at each predicate’s ending time, the feasible sets of all previously active HOCBFs remain mutually compatible. To illustrate the idea, consider two predicates $\mu_{1}$ and $\mu_{2}$ with ending times $t_{b}^{1}\leq t_{b}^{2}$. Since $b_{1}$ must be maintained until $t_{b}^{1}$, we require that at $t_{b}^{1}$ the feasible region induced by $b_{2}$ does not conflict with that of $b_{1}$, then the corresponding compatibility condition becomes:
@@
Because $\gamma(t,\bm{\omega})$ in (21) and (23) is non-increasing, the temporal operators are enforced as follows:
@@
  • For $F_{[t_{a},t_{b}]}(\bigwedge_{j=1}^{N}\mu_{j})$, constraint (22b) ensures the existence of a common $t^{\prime}\in[t_{a},t_{b}]$ where all $h_{j}(\mathbf{x}(t^{\prime}))\geq 0$.
@@
Consider system (1) satisfying Assumptions 1 and 2, and HOCBFs constructed by (19) and (20) that satisfy (22), (24), and (27). Suppose each HOCBF satisfies $\psi_{i,j}(\mathbf{x}_{0},0)\geq 0$ for $i=0,\ldots,m-1$, $j=1,\ldots,M$. Then any control input $\mathbf{u}(t)$ that satisfies (7) for all relevant HOCBFs guarantees satisfaction of the STL specification $\varphi$.
@@
Constraints (22a) and (24a) ensure $b_{j}(\mathbf{x}_{0},0)\geq 0$ for all predicates. Since $\psi_{i,j}(\mathbf{x}_{0},0,\boldsymbol{\theta}_{b})\geq 0$ for all $i$, Theorem 1 guarantees that any control input satisfying (7) yields $b_{j}(\mathbf{x},t)\geq 0$ for all undeleted HOCBFs. Because $\gamma_{j}(t,\bm{\omega}_{j})$ is non-increasing, (22b) ensures $\exists t^{\prime}\in[t_{a},t_{b}]$, $h_{j}(\mathbf{x}(t^{\prime}))\geq 0$ for $F{[t_{a},t_{b}]}\mu$, while (24b) ensures $\forall t^{\prime}\in[t_{a},t_{b}]$, $h_{j}(\mathbf{x}(t^{\prime}))\geq 0$ for $G{[t_{a},t_{b}]}\mu$. For a formula in the form of $F_{[t_{a},t_{b}]}(\bigwedge_{j=1}^{N}\mu_{j})$, (22b) ensures $\exists t^{\prime}\in[t_{a},t_{b}]$, $h_{j}(\mathbf{x}(t^{\prime}))\geq 0$ for all $j$. Hence every predicate in $\varphi$ is satisfied over its required temporal interval, ensuring $(\mathbf{x},0)\models\varphi$. ∎
@@ IV-C Unified Robustness Metric
In Sec. IV-B we constructed either (i) a time-invariant HOCBF $b_{j}(\mathbf{x})$ (for Category I predicates), or (ii) a time-varying HOCBF $b_{j}(\mathbf{x},t,\bm{\theta}_{b})$ (for Category II and III predicates), together with a deletion rule that removes the corresponding constraint once the predicate is satisfied (for $F_{[t_{a},t_{b}]}\mu$ operators) or at the terminal time (for $G_{[t_{a},t_{b}]}\mu$ operators). Before deletion, each HOCBF imposes a highest-order constraint $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})\geq 0$, which must remain feasible for the dQP (10). This subsection introduces a unified robustness combining: (i) feasibility robustness, and (ii) STL temporal robustness based on the exponential robustness in [21], ensuring that the controller remains feasible while satisfying the STL specification.
@@
Feasibility robustness. As shown in Theorem 3, the dQP is feasible at time $t$ whenever $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$. Thus the instantaneous feasibility margin is
$\rho^{\mathrm{fea}}_{j}(\mu_{j},\mathbf{x},t):=\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p}).$ (31)
This margin must stay positive until the HOCBF is deleted. An HOCBF corresponding to a predicate $\mu_{j}$ enclosed by $F_{[t_{a}^{j},t_{b}^{j}]}$ is deleted once $h_{j}(\mathbf{x})\geq 0$. We use an additional STL subformula to express the feasibility requirement:
$\varphi_{j}^{\mathrm{fea}}:=\big(\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0\big)U_{[t_{a}^{j},t_{b}^{j}]}\big(h_{j}(\mathbf{x})\geq 0\big),$ (32)
where $U_{[t_{a}^{j},t_{b}^{j}]}$ is the temporal Until operator as defined in standard STL. An HOCBF corresponding to a predicate $\mu_{j}$ enclosed by $G_{[t_{a}^{j},t_{b}^{j}]}$ is always deleted at $t_{b}^{j}$. Similiarly, we define the feasibility STL subformula as:
$\varphi_{j}^{\mathrm{fea}}:=G_{[t_{a}^{j},t_{b}^{j}]}\big(\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0\big).$ (33)
With the construction above, we now have $M$ predicates $\mu_{j}$, $j=1,\ldots,M$, spanning Categories I–III. Each predicate $\mu_{j}$ is associated with an STL subformula $\varphi_{j}^{\mathrm{fea}}$ that enforces QP feasibility until deletion, given by expression (32) and (33).
To capture the control bounds (10c), we introduce another STL formula defined over the control signal $\mathbf{u}$ which must hold for all $t\in[0,T]$:
$\varphi^{\mathcal{U}}:=G_{[0,T]}\big((\mathbf{u}(t)-\mathbf{u}_{\min}\geq\mathbf{0})\wedge(\mathbf{u}_{\max}-\mathbf{u}(t)\geq\mathbf{0})\big).$ (34)
Note that $\varphi_{j}^{fea}$ and $\varphi^{\mathcal{U}}$ in (32), (33) and (34) are additional STL formulas other than the original STL specification defined in (4), so they have different syntax.
@@
The robustness of $\widetilde{\varphi}$ can be evaluated using the exponential robustness in (28)–(30). We denote the resulting unified scalar robustness by
@@
Theorem 5 (Feasibility-Aware Correctness of BarrierNet Controllers).
Consider the system (1) and the STL specification $\varphi$ satisfying Assumptions 1 and 2. Construct the HOCBFs and their feasibility subformulas $\varphi_{j}^{\mathrm{fea}}$ as in (32), (33), and define the augmented STL specification $\widetilde{\varphi}$ by combining $\varphi$ with all feasibility and control-bound subformulas as in (35). Let a neural network controller be given whose last layer is the differentiable QP (10), which enforces all highest-order HOCBF constraints, deletion rules, and control bounds at all time. If the dQP (10) remains feasible for all $t>0$, then the closed-loop trajectory generated by this controller guarantees the satisfaction of the original STL specification $\varphi$. Moreover, if each STL subformula $\varphi_{j}^{\mathrm{fea}}$ is satisfied, i.e., $(\mathbf{x},0)\models\bigwedge_{j=1}^{M}\varphi_{j}^{\mathrm{fea}}$, then the dQP (10) remains feasible for all $t\geq 0$.
@@
By Theorems 2 and 4, the controller satisfying (10) guarantees the resulting closed-loop trajectory satisfies the STL specification $\varphi$. By satisfaction of each feasibility subformula $\varphi_{j}^{\mathrm{fea}}$, we have that $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$ for all $t$ prior to deletion. By Theorem 3, the dQP remains pointwise feasible at all times. ∎
Corollary 1.
If the conditions of Theorem 5 hold, then $\rho^{\mathrm{uni}}(\mathbf{x},0):=\rho(\widetilde{\varphi},\mathbf{x},0)\geq 0$.
Proof.
By Theorem 5, the stated controller conditions imply $(\mathbf{x},0)\models\widetilde{\varphi}$. By soundness of exponential robustness, $(\mathbf{x},0)\models\widetilde{\varphi}$ implies $\rho(\widetilde{\varphi},\mathbf{x},0)\geq 0$, i.e., $\rho^{\mathrm{uni}}(\mathbf{x},0)\geq 0$. ∎
Remark 2.
The original STL specification $\varphi$ is theoretically guaranteed in continuous time through HOCBF constraints embedded in the dQP formulation. However, in practice, the resulting QP is implemented in discrete time, which introduces inter-sampling effects [20] and may affect both the continuous-time satisfaction of the specification and its discrete-time evaluation at the sampling instants. Therefore, the robustness of the original STL specification $\varphi$ in (35) is monitored to reflect its satisfaction under discrete-time enforcement. The feasibility-related STL subformulas $\bigwedge_{j=1}^{M}\varphi_{j}^{\mathrm{fea}}$ are not enforced by hard constraints and thus are not strictly guaranteed. Instead, their robustness values, together with the satisfaction of the input-bound specification $\varphi^{\mathcal{U}}$, are incorporated as optimization objectives during training to guide and improve feasibility. Addressing inter-sampling effects is left for future work.